OpenStudy (anonymous):

A glass vase is shaped like a cylinder and has a radius of 2.5 inches. The vase holds 170 cubic inches of water. How tall is the vase?

OpenStudy (anonymous):

@wmj259

OpenStudy (isaiah.feynman):

Volume of a cylinder is \[V = \pi r^{2}h \]

OpenStudy (anonymous):

it's not asking the volume, its asking the height though.. that means i divide right?

OpenStudy (isaiah.feynman):

We are looking for h. So we can solve for h and then substitute known values. \[h = \frac{ V }{ \pi r^{2} }\]

OpenStudy (isaiah.feynman):

What is V and r?

OpenStudy (anonymous):

v= 170 r= 2.5